# Rate-Distortion-Perception Tradeoff for Gaussian Vector Sources

### v1

@@ IV-D RDP Function and Generalized Reverse Water-filling with Wasserstein-2 Distance as Perception Metric
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}(D,P)}{\hat{\lambda}^{*}_{{\ell}}(D,P)}}+1\right)+\nu_{2}\left(1-\sqrt{\frac{\lambda_{\ell}}{\hat{\lambda}^{*}_{\ell}(D,P)}}\right)+\eta_{\ell}=0.$ (50)
@@
Let $(D,P)$ be a given distortion and perception constraints that are strictly feasible. The optimal solution of (34) with the perception metric (33) replaced by (48) is given as follows:
@@
    $\left\{\hat{\lambda}_{\ell}\ \left|\ \sum_{\ell=1}^{L}{P}_{\ell}(\hat{\lambda}_{\ell})\leq P,\ \ \sum_{\ell=1}^{L}\lambda_{\ell}+\hat{\lambda}_{\ell}\leq D,\ \ \hat{\lambda}_{\ell}\geq 0\right.\right\}.$ (57)
@@ IV-E Perceptually Perfect Reconstruction
$\gamma^{*}_{\ell}(D,0)=\frac{2\lambda_{\ell}}{1+\sqrt{1+16\nu_{1}^{2}\lambda_{\ell}^{2}}},\qquad\ell\in\{1,\ldots,L\}.$ (60)
@@
An interpretation of the optimal rate allocation in this $P=0$ case is as follows. By (58), the optimal rate allocated to the $\ell$-th component is controlled by the expression $\frac{1+\sqrt{1+16\nu_{1}^{2}\lambda_{\ell}^{2}}}{2}$. So, if a component has a larger variance, it is compressed at a higher rate. Further, by (60) it also has a higher water-level.
@@
where the water-levels are given by
$\gamma^{*}_{\ell}\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)=\lambda_{\ell}-\frac{\epsilon^{2}\lambda_{\ell}^{3}}{4\left(\sum_{\ell=1}^{L}\lambda_{\ell}^{2}\right)^{2}}+O(\epsilon^{3}),\quad\ell\in\{1,\ldots,L\}.$ (62)
@@
See Appendix H-1.∎
Here, we express $R(D,0)$ in term of deviation from the maximum distortion at perfect perception at zero rate. This maximum distortion can be shown to be $2\sum_{\ell=1}^{L}\lambda_{\ell}$, which is twice of the total variance of the source [9], because at zero rate the decoder should simply generate an independent Gaussian random vector with the same covariance matrix. Comparing $R\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)$ of Proposition 3 with $R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ in Proposition 1, it is interesting to see that the variances of the source enter $R\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)$ as $\sum_{\ell=1}^{L}\lambda^{2}_{\ell}$ which is the sum of the square of the variances over all the components. This is in contrast to the corresponding factor in $R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ in the traditional reverse water-filling solution which is simply $\lambda^{\max}$. This is a consequence of the perfect perception constraint, which requires all the components to be reconstructed with the same variances as the source at the decoder.
@@
$R(\epsilon,0)=\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}+\frac{\epsilon}{8L}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{2}),$ (63)
where the water-levels are given by
$\gamma^{*}_{\ell}(\epsilon,0)=\frac{\epsilon}{L}-\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}+\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),\quad\ell\in\{1,\ldots,L\}.$ (64)
@@
$\gamma^{*}_{\ell}(\epsilon,\infty)-\gamma^{*}_{\ell}(\epsilon,0)$ $=$ $\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}-\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),\quad\ell\in\{1,\ldots,L\}.$ (66)
@@
Figure 3: The water-levels assigned to different components for a Gaussian vector source with $\lambda_{1}=3,\lambda_{2}=2,\lambda_{3}=5,\lambda_{4}=4$ and $\lambda_{5}=1$.
Fig. 3 shows the water-levels of different components for both low-distortion and high-distortion compression with $P=\infty$ or $P=0$ for an example of a Gaussian vector source. The water-levels determine the compression rates assigned to each component.
In Fig. 3(a), for high-distortion compression with no perception constraint, all components except the one with the largest eigenvalue are allocated a zero compression rate (cf. Proposition 1). With an active perception constarint, as shown in Fig. 3(c) for the $P=0$ case, all components are allocated positive rates (cf. Proposition 3).
In Fig. 3(b), for low-distortion compression with no perception constraint, the water-levels of all components are the same (cf. Proposition 2). At low distortion and with an active perception constraint, as shown in Fig. 3(d) for the $P=0$ case, the water-levels of different components are approximately equal with some slight differences which are determined by (64) in Proposition 4. Therefore, in the low-distortion regime, the water-levels of all components are approximately the same regardless of the perception constraint.
@@ V Conclusions
This paper characterizes the RDP function for a Gaussian vector source. In contrast to the traditional reverse water-filling solution (without a perception constraint), the water-levels assigned to different components are not necessarily equal. When both distortion and perception constraints are active, every component is assigned a positive rate. These results have implications to perception-aware image coding.
@@ Appendix A Asymptotic Analysis of the Traditional RD Function
A-1 High-Distortion Compression
@@
First consider the case that $|L^{\max}|=1$. The distortion constraint (17) implies that
@@
If $|L^{\max}|>1$, then similar to the above discussion, all eigenvalues except the largest ones are assigned a zero compression rate and for the maximum eigenvalues, we have the following water-level
$\nu(D)=\lambda^{\max}-\frac{\epsilon}{|L^{\max}|},$ (73)
@@
$R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\frac{|L^{\max}|}{2}\log\frac{\lambda_{L}}{\lambda_{L}-\frac{\epsilon}{|L^{\max}|}}$ (74)
@@
This proves (22) for arbitrary $L^{\max}$.
A-2 Low-Distortion Compression
Consider the case of $D=\epsilon$ for sufficiently small $\epsilon>0$. In this low-distortion regime, the constant water-level $\nu(D)$ is not saturated by the eigenvalues. Thus, Proposition 1 simplifies to the following
@@ Appendix B Proof of Theorem 2
$=$ $-h(\hat{X}^{*})+\frac{1}{2}\int P_{\hat{X}^{*}}(x)x\Sigma_{X}^{-1}x^{T}dx+\frac{1}{2}\log(2\pi)^{L}\det(\Sigma_{X})$ (86)
$=$ $-h(\hat{X}^{*})+\frac{1}{2}\int P_{\hat{X}_{G}}(x)x\Sigma_{X}^{-1}x^{T}dx+\frac{1}{2}\log(2\pi)^{L}\det(\Sigma_{X})$ (87)
@@
Thus, replacing $\hat{X}^{*}$ by $\hat{X}_{G}$ does not increase the rate, while distortion and perception constraints remain to be satisfied. Thus, the optimal $\hat{X}^{*}$ must be jointly Gaussian with $X$.
@@
Now, define $P_{U_{G}V_{G}}$ to be a joint Gaussian distribution such that
@@
$P\geq W_{2}^{2}(P_{X},P_{\hat{X}^{*}})$ $=$ $\inf_{\begin{subarray}{c}\tilde{P}_{UV}:\\ \tilde{P}_{U}=P_{X}\\ \tilde{P}_{V}=P_{\hat{X}^{*}}\end{subarray}}\mathbbm{E}_{\tilde{P}}[\|U-V\|^{2}]$ (93)
@@
Thus, without loss of optimality one can replace $\hat{X}^{*}$ by $\hat{X}_{G}$ since the rate does not increase, while the distortion and perception constraints remain to be satisfied.
@@ Appendix C Proof of Theorem 3
$R^{*}(D,P)=\min_{\{\hat{\lambda}_{\ell},\gamma_{{\ell}}\}_{\ell=1}^{L}}\;\;\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}$ (101a)
@@
$\hskip 113.81102pt0\leq\hat{\lambda}_{\ell},$ (101c)
$\hskip 113.81102pt\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}}\right)\leq D,$ (101d)
$\hskip 113.81102pt\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)\leq P.$ (101e)
C-1 Proof of $R^{*}(D,P)\geq R(D,P)$
@@ C-2 Proof of $R^{*}(D,P)\leq R(D,P)$
$\gamma_{\ell}=\mathbbm{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}],\qquad\ell\in\{1,\ldots,L\}.$ (116)
@@
$\geq$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}|\hat{Z}^{*}_{G,\ell})$ (120)
$=$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}]|\hat{Z}^{*}_{G,\ell})$ (121)
$=$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])$ (122)
$=$ $\sum_{\ell=1}^{L}\frac{1}{2}\log\left((2\pi e)\lambda_{{\ell}}\right)-\sum_{\ell=1}^{L}\frac{1}{2}\log\left((2\pi e)\gamma_{{\ell}}\right)$ (123)
@@
  • (117) is due to the invertibility of unitary transformations,
  • (119) follows because $Z_{1},\ldots,Z_{L}$ are independent,
  • (120) follows from the chain rule and that conditioning does not increase entropy,
  • (122) follows because $Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}]$ is independent of $\hat{Z}^{*}_{G,\ell}$,
  • (123) follows because $\mathbb{E}[Z_{\ell}^{2}]=\lambda_{{\ell}}$ and $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]=\gamma_{{\ell}}$.
@@
$=$ $\sum_{\ell=1}^{L}\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}}$ (129)
@@
  • (129) follows from the identity $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]=\mathbb{E}[Z_{\ell}^{2}]-(\mathbb{E}[Z_{\ell}\hat{Z}^{*}_{G,\ell}])^{2}(\mathbb{E}[\hat{Z}^{*}_{G,\ell}])^{-1}$, and $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]=\gamma_{\ell}$, $\mathbb{E}[Z_{\ell}^{2}]=\lambda_{{\ell}}$, $\mathbb{E}[(\hat{Z}^{*}_{G,\ell})^{2}]=\hat{\lambda}_{{\ell}}$.
@@
$P\geq D(P_{\hat{X}^{*}_{G}}\|P_{X})$ $=$ $\frac{1}{2}\left(\text{tr}(\Lambda_{X}^{-1}\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T})-L+\log\frac{\det(\Lambda_{X})}{\det(\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T})}\right)$ (130)
@@
  • (132) follows from Hadamard’s inequality for a positive semidefinite matrix.

### v2

@@ IV-D RDP Function and Generalized Reverse Water-filling with Wasserstein-2 Distance as Perception Metric
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}(D,P)}{\hat{\lambda}^{*}_{{\ell}}(D,P)}}+1\right)$
$\hskip 28.45274pt+\nu_{2}\left(1-\sqrt{\frac{\lambda_{\ell}}{\hat{\lambda}^{*}_{\ell}(D,P)}}\right)+\eta_{\ell}=0.$ (53)
@@
Let $(D,P)$ be given distortion and perception constraints that are strictly feasible. The optimal solution of (36) with the perception metric (35) replaced by (51) is given as follows:
@@
    $\Bigg{\{}\hat{\lambda}_{\ell}\Bigg{|}\sum_{\ell=1}^{L}{P}_{\ell}(\hat{\lambda}_{\ell})\leq P,\sum_{\ell=1}^{L}\lambda_{\ell}+\hat{\lambda}_{\ell}\leq D,\hat{\lambda}_{\ell}\geq 0\Bigg{\}}.$ (60)
@@ IV-E Perceptually Perfect Reconstruction
$\gamma^{*}_{\ell}(D,0)=\frac{2\lambda_{\ell}}{1+\sqrt{1+16\nu_{1}^{2}\lambda_{\ell}^{2}}},\;\;\;\ell\in\{1,\ldots,L\}.$ (63)
@@
An interpretation of the optimal rate allocation in this $P=0$ case is as follows. By (61), the optimal rate allocated to the $\ell$-th component is controlled by the expression $\frac{1+\sqrt{1+16\nu_{1}^{2}\lambda_{\ell}^{2}}}{2}$. So, if a component has a larger variance, it is compressed at a higher rate. Further, by (63) it also has a higher water level.
@@
where the water levels are given by
$\gamma^{*}_{\ell}\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)$ $=$ $\lambda_{\ell}-\frac{\epsilon^{2}\lambda_{\ell}^{3}}{4\left(\sum_{\ell=1}^{L}\lambda_{\ell}^{2}\right)^{2}}+O(\epsilon^{3}),$ (65)
$\hskip 56.9055pt\ell\in\{1,\ldots,L\}.$
@@
See Appendix H-A.∎
Here, we express $R(D,0)$ in term of deviation from the maximum distortion at perfect perception at zero rate. This maximum distortion can be shown to be $2\sum_{\ell=1}^{L}\lambda_{\ell}$, which is twice of the total variance of the source [9], because at zero rate the decoder should simply generate an independent Gaussian random vector with the same covariance matrix. Comparing $R\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)$ in Proposition 3 with $R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ in Proposition 1, it is interesting to see that the variances of the source appear in $R\left(2\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,0\right)$ as $\sum_{\ell=1}^{L}\lambda^{2}_{\ell}$ which is the sum of the square of the variances over all the components. This is in contrast to the corresponding factor in $R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ in the traditional reverse water-filling solution which is simply $\lambda^{\max}$. This is a consequence of the perfect perception constraint, which requires all the components to be reconstructed with the same variances as the source at the decoder.
@@
$R(\epsilon,0)=\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{L\lambda_{\ell}}{\epsilon}+\frac{\epsilon}{8L}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{2}),$
where the water levels are given by
$\gamma^{*}_{\ell}(\epsilon,0)$ $=$ $\frac{\epsilon}{L}-\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}+\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),$ (67)
$\hskip 99.58464pt\ell\in\{1,\ldots,L\}.$
@@
$\gamma^{*}_{\ell}(\epsilon,\infty)-\gamma^{*}_{\ell}(\epsilon,0)=\frac{\epsilon^{2}}{2L^{2}\lambda_{\ell}}-\frac{\epsilon^{2}}{4L^{3}}\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}+O(\epsilon^{3}),$
$\hskip 150.79968pt\ell\in\{1,\ldots,L\}.$ (69)
@@
Figure 3: The water levels assigned to different components for a Gaussian vector source with $\lambda_{1}=3,\lambda_{2}=2,\lambda_{3}=5,\lambda_{4}=4$ and $\lambda_{5}=1$.
Fig. 3 shows the water levels of different components for both low-distortion and high-distortion compression with $P=\infty$ or $P=0$ for an example of a Gaussian vector source. The water levels determine the compression rates assigned to each component.
In Fig. 3(a), for high-distortion compression with no perception constraint, all components except the one with the largest eigenvalue are allocated a zero compression rate (cf. Proposition 1). With an active perception constraint, as shown in Fig. 3(c) for the $P=0$ case, all components are allocated positive rates (cf. Proposition 3).
In Fig. 3(b), for low-distortion compression with no perception constraint, the water levels of all components are the same (cf. Proposition 2). At low distortion and with an active perception constraint, as shown in Fig. 3(d) for the $P=0$ case, the water levels of different components are approximately equal with some slight differences which are determined by (67) in Proposition 4. Therefore, in the low-distortion regime, the water levels of all components are approximately the same regardless of the perception constraint.
@@ V Conclusions
This paper characterizes the RDP function for a Gaussian vector source. In contrast to the traditional reverse water-filling solution (without a perception constraint), the water levels assigned to different components are not necessarily equal. When both distortion and perception constraints are active, every component is assigned a positive rate. These results have implications to perception-aware image coding.
@@ Appendix A Asymptotic Analysis of the Traditional RD Function
A-A High-Distortion Compression
@@
First consider the case that $|\mathcal{L}^{\max}|=1$. The distortion constraint (17) implies that
@@
If $|\mathcal{L}^{\max}|>1$, then similar to the above discussion, all eigenvalues except the largest ones are assigned a zero compression rate and for the maximum eigenvalues, we have the following water level
$\nu(D)=\lambda^{\max}-\frac{\epsilon}{|\mathcal{L}^{\max}|},$ (76)
@@
$R\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\frac{|\mathcal{L}^{\max}|}{2}\log\frac{\lambda_{L}}{\lambda_{L}-\frac{\epsilon}{|\mathcal{L}^{\max}|}}$ (77)
@@
This proves (22) for arbitrary $\mathcal{L}^{\max}$.
A-B Low-Distortion Compression
Consider the case of $D=\epsilon$ for sufficiently small $\epsilon>0$. In this low-distortion regime, the constant water level $\nu(D)$ is not saturated by the eigenvalues. Thus, Proposition 1 simplifies to the following
@@ Appendix B Proof of Theorem 2
[…]
$\hskip 14.22636pt+\frac{1}{2}\log(2\pi)^{L}\det(\Sigma_{X})$
[…]
$\hskip 14.22636pt+\frac{1}{2}\log(2\pi)^{L}\det(\Sigma_{X})$
@@
Thus, replacing $\hat{X}^{*}$ by $\hat{X}_{G}$ does not increase the rate, while distortion and perception constraints remain satisfied. Thus, the optimal $\hat{X}^{*}$ must be jointly Gaussian with $X$.
@@
Let $P_{U_{G}V_{G}}$ be a joint Gaussian distribution such that
@@
$P$$\geq$ $W_{2}^{2}(P_{X},P_{\hat{X}^{*}})$ (96)
$=$ $\inf_{\begin{subarray}{c}\tilde{P}_{UV}:\\ \tilde{P}_{U}=P_{X}\\ \tilde{P}_{V}=P_{\hat{X}^{*}}\end{subarray}}\mathbbm{E}_{\tilde{P}}[\|U-V\|^{2}]$ (97)
@@
Thus, without loss of optimality one can replace $\hat{X}^{*}$ by $\hat{X}_{G}$ since the rate does not increase, while the distortion and perception constraints remain satisfied.
@@ Appendix C Proof of Theorem 3
$R^{*}(D,P)$
$=\min_{\{\hat{\lambda}_{\ell},\gamma_{{\ell}}\}_{\ell=1}^{L}}\frac{1}{2}\sum_{\ell=1}^{L}\log\frac{\lambda_{\ell}}{\gamma_{{\ell}}}$ (105a)
@@
$\hskip 51.21504pt0\leq\hat{\lambda}_{\ell}$ (105c)
$\hskip 51.21504pt\sum_{\ell=1}^{L}\left(\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}}\right)\leq D\;\;\;\;\;\;$ (105d)
$\hskip 51.21504pt\frac{1}{2}\sum_{\ell=1}^{L}\left(\frac{\hat{\lambda}_{{\ell}}}{\lambda_{{\ell}}}-1+\log\frac{\lambda_{{\ell}}}{\hat{\lambda}_{{\ell}}}\right)\leq P.$ (105e)
C-A Proof of $R^{*}(D,P)\geq R(D,P)$
@@ C-B Proof of $R^{*}(D,P)\leq R(D,P)$
$\gamma_{\ell}=\mathbbm{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}],\;\;\;\ell\in\{1,\ldots,L\}.$ (120)
@@
$\geq$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}|\hat{Z}^{*}_{G,\ell})$
$=$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}]|\hat{Z}^{*}_{G,\ell})$
$=$ $\sum_{\ell=1}^{L}h(Z_{\ell})-\sum_{\ell=1}^{L}h(Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])$
$=$ $\sum_{\ell=1}^{L}\frac{1}{2}\log\left((2\pi e)\lambda_{{\ell}}\right)-\sum_{\ell=1}^{L}\frac{1}{2}\log\left((2\pi e)\gamma_{{\ell}}\right)$
@@
  • (121) is due to the invertibility of unitary transformations;
  • (123) follows because $Z_{1},\ldots,Z_{L}$ are independent;
  • (C-B) follows from the chain rule and that conditioning does not increase entropy;
  • (C-B) follows because $Z_{\ell}-\mathbbm{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}]$ is independent of $\hat{Z}^{*}_{G,\ell}$;
  • (LABEL:eq4) follows because $\mathbb{E}[Z_{\ell}^{2}]=\lambda_{{\ell}}$ and $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]=\gamma_{{\ell}}$.
@@
$=$ $\sum_{\ell=1}^{L}\lambda_{{\ell}}-2\sqrt{\hat{\lambda}_{{\ell}}(\lambda_{{\ell}}-\gamma_{{\ell}})}+\hat{\lambda}_{{\ell}},$ (134)
@@
  • (134) follows from the identity
    $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]=\mathbb{E}[Z_{\ell}^{2}]-(\mathbb{E}[Z_{\ell}\hat{Z}^{*}_{G,\ell}])^{2}(\mathbb{E}[\hat{Z}^{*}_{G,\ell}])^{-1}$ (135)
    and
    $\mathbb{E}[(Z_{\ell}-\mathbb{E}[Z_{\ell}|\hat{Z}^{*}_{G,\ell}])^{2}]$ $=$ $\gamma_{\ell},$ (136)
    $\mathbb{E}[Z_{\ell}^{2}]$ $=$ $\lambda_{{\ell}},$ (137)
    $\mathbb{E}[(\hat{Z}^{*}_{G,\ell})^{2}]$ $=$ $\hat{\lambda}_{{\ell}}.$ (138)
@@
$P$$\geq$ $D(P_{\hat{X}^{*}_{G}}\|P_{X})$
$=$ $\frac{1}{2}\Bigg{(}\text{tr}(\Lambda_{X}^{-1}\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T})-L+\log\frac{\det(\Lambda_{X})}{\det(\Theta\Sigma_{\hat{X}^{*}_{G}}\Theta^{T})}\Bigg{)}$
@@
  • (142) follows from the Hadamard’s inequality for a positive semidefinite matrix.